OpenStudy (anonymous):

A rectangle has a perimeter of 80 cm. If its width is x, express its length and its area in terms of x. What is the maximum area of the rectangle?

OpenStudy (ybarrap):

2x + 2l = 80 2l = 80 - 2x l = (80 - 2x)/2 = 40 - x A = Area = l*x = (40 - x)*x = 40x - x^2 dA/dx = 40 - 2x dA/dx = 0 = 40 - 2x 2x = 40 x = 40/2 = 20 l = 40 - x = 40 - 20 = 20 max area = 20*20 = 400 cm^2 Ans. 400 cm^2